OpenStudy (<3pandas18):

Find the values of the variables and the lengths of the sides of this kite.

OpenStudy (<3pandas18):

Plse help ASAP

OpenStudy (anonymous):

In a kite, the top two sides are equal to each other and the bottom two sides are also equal to each other. So you cant set y-4 = x+5 and 2x+5 = x + 12

OpenStudy (anonymous):

oops typo. So you *can* set

OpenStudy (anonymous):

Solve for x first, since that's the variable present in both equations, and use it to solve for y. So 2x + 5 = x + 12 2x - x = 12 - 5 (x = 7) <-- plug 7 into x in the first equation (y-4 = x+5) ------------------------------------------------ and you get: y-4 = 7 + 5 solve for y... (y = 16)

OpenStudy (anonymous):

& now you can find the length of the sides too :)

OpenStudy (<3pandas18):

THANK YOU!!!!!